### v3

# Exploring lattice thermal conductivity models via interpretable deep learning to accelerate the discovery of novel materials

Abstract
Lattice thermal conductivity, being integral to thermal transport properties, is indispensable to advancements in areas such as thermoelectric materials and thermal management. Traditional methods, such as Density Functional Theory and Molecular Dynamics, require significant computational resources, posing challenges to the high-throughput prediction of lattice thermal conductivity. Although AI-driven material science has achieved fruitful progress, the trade-off between accuracy and interpretability in machine learning continues to hinder further advancements. This study utilizes interpretable deep learning techniques to construct a rapid prediction framework that enables both qualitative assessments and quantitative predictions, accurately forecasting the thermal transport properties of three novel materials. Furthermore, interpretable deep learning offers analytically grounded physical models while integrating with sensitivity analysis to uncover deeper theoretical insights.
I Introduction
Lattice Thermal Conductivity (LTC) is a critical physical parameter that quantifies a material’s ability to transfer heat through lattice vibrations. It has diverse applications in thermal management[1], energy conversion[2], and thermoelectric materials[3]. High-LTC materials are used in electronic devices to efficiently dissipate heat and prevent overheating[4]. In contrast, low-LTC materials exhibit excellent performance in thermoelectric conversion, making them ideal for developing efficient thermoelectric generators (TEGs)[5] and coolers[6]. LTC is crucial for material design and optimization, yet acquiring it for specific materials is challenging.
Traditional experimental methods for measuring LTC, such as the laser flash method[7] and thermal conductivity probes[8], are often inefficient. In theoretical calculations, solving the Boltzmann Transport Equation (BTE) based on Density Functional Theory (DFT) is regarded as the most reliable method for determining LTC[9, 10]. Molecular Dynamics (MD) offers an alternative[11]; however, the former is limited by its substantial computational resource requirements, while the latter’s accuracy depends on the choice of interatomic potentials[12]. In recent years, machine learning (ML) has emerged as a powerful and efficient data mining tool, gaining widespread application in materials science[13]. The intersection of materials science and artificial intelligence is commonly referred to as “Materials Informatics”[14] or the “Materials Genome”[15]. Early ML-based predictions of material properties primarily aimed at achieving high prediction accuracy[16, 17, 18]. Nevertheless, such efforts, relying on “black-box” models, offered limited support for advancing theoretical research in materials science. The focus has shifted towards model interpretability, prompting greater adoption of “white-box” models that contribute more substantially to theoretical advancements[19]. Research in materials informatics related to thermal conductivity has followed this trend. Efforts to predict LTC using black-box models such as Random Forest (RF)[20], Gaussian Process Regression (GPR)[21], and eXtreme Gradient Boosting (XGBoost)[22] have yielded reliable accuracy, whereas, the interpretability of these models is often hindered by their complexity. Alternatively, Genetic Programming-based Symbolic Regression (GPSR)[23] algorithms offer better interpretability, but their simplicity comes at the cost of reduced accuracy. The prevailing view is that the behavior and processes of complex models are difficult to understand and interpret, while simple models often lack strong fitting capabilities[24, 25, 19]. This presents a challenging trade-off, but the situation is gradually improving. The Sure Independence Screening and Sparsifying Operator (SISSO)[26], based on compressed sensing and symbolic regression, has not only surpassed the accuracy of the Slack semi-empirical model[27] but also narrowed the gap with black-box models like Kernel Ridge Regression (KRR) and GPR, and has been used to quantify feature sensitivity and identify key physical parameters influencing LTC[28]. Despite its potential, SISSO’s applicability is constrained by its high computational demands. Enumerating combinations of features and operators to construct descriptors poses an NP-hard problem[26], with resource requirements increasing for high-dimensional input features[29].
Figure 1: A schematic illustration of the DL framework for LTC modeling and high-throughput prediction. This framework facilitates accurate and interpretable predictions of LTC for inorganic crystalline materials.
Kolmogorov-Arnold Networks (KANs)[30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving partial differential equations (PDEs), time series prediction, and high-dimensional data tasks[31, 30, 32]. In some cases, smaller KAN models outperform MLPs[31, 30]. KANs have also demonstrated success in engineering applications, such as predicting chiller energy consumption[33] and nuclear binding energy[34], highlighting both feasibility and broad potential. Notably, their capabilities in symbolic regression offer interpretability advantages tabsent in MLPs[31, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict LTC. However, achieving accurate LTC predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
Figure 1 illustrates the framework of our work, which aims to model and predict LTC using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to LTC. This approach further elucidates the relationships between LTC and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN)[35] is employed to predict critical physical features required for LTC modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Finally, by integrating DFT and MD, our pretrained KAN model achieves outstanding accuracy in predicting the properties of candidates in an unlabeled extrapolation dataset.
II Results
II.1 Establishment and evaluation of LTC models
Table 1: The ranges for hyperparameter optimization.
Model Hyperparameter Range/Value
MLP hidden_sizes [64, 64, 64] (fixed)
learning_rate [0.001, 0.1]
batch_size [64, 256]
num_epochs [100, 600]
XGB n_estimators [50, 500]
learning_rate [0.0001, 0.5]
max_depth [1, 6]
subsample [0.8, 1]
colsample_bytree [0.8, 1]
reg_alpha [0.8, 1]
reg_lambda [5, 50]
KAN width [5, 15]
grid [5, 15]
k [2, 10]
lamb_l1 [5, 50]
steps [20, 35]
lr [0.01, 1.0]
In selecting black-box models, we consider XGBoost and MLP as our primary options. XGBoost, a representative of ensemble learning, has demonstrated notable performance in various applications, including electrocaloric temperature change prediction in ceramics[17] and materials mechanical property prediction[36]. XGBoost is an enhancement of gradient boosting decision trees (GBDT). Compared to traditional GBDT, XGBoost introduces several innovations, including regularization (to improve generalization), a two-step gradient approximation for the objective function (to enhance computational efficiency), column subsampling (to reduce noise and boost generalization), and handling missing values (using a default direction for tree nodes, making it suitable for sparse datasets)[102]. MLP is a type of feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer. Each neuron in a given layer is fully connected to all neurons in the preceding layer through weight matrices, allowing the network to perform linear combinations and nonlinear mappings of high-dimensional features. Neurons in each hidden layer typically use nonlinear activation functions such as ReLU, which endow the MLP with considerable expressive power, enabling it to approximate arbitrarily complex nonlinear functions. The training process of an MLP is carried out using the backpropagation algorithm, which optimizes the network’s weights through gradient descent to minimize a loss function. While MLPs exhibit substantial performance advantages in handling complex data patterns and feature learning tasks, their computational complexity and sensitivity to hyperparameter choices can make the training process time-consuming[38].
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch[39] and scikit-learn[40] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna[41] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal[42]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in Table S1.
In contrast to the previously discussed algorithms, SISSO imposes the highest computational demands during model training. Therefore, we followed hyperparameter configurations recommended in established methodologies from related work[28]. To enhance interpretability, we maintained strict dimensional consistency across all features in SISSO, effectively eliminating any potential for invalid operations.
Figure 2: Model performance on dataset with complete features. (a) The blue circular markers represent samples from the training set, while the red star scatters are samples from the test set. (b) KAN model architecture based on the original dataset.
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria[43]. As shown in Figure 2(a), on the original dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low LTC values (approximately $log(\kappa)<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling[44] and weighting[44, 45]. However, these methods inevitably impact the model’s interpretability[46, 47]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
  • Molecular features are often non-smooth, whereas deep models tend to favor smooth solutions[48, 49];
  • Different dimensions of molecular features typically carry distinct information, yet deep models tend to integrate features across dimensions[49];
  • Tree-based models like XGBoost inherently handle redundant features, while deep models are more susceptible to interference from such features[48];
  • MLP is rotation-invariant, and any rotation-invariant learning process inherently exhibits the worst-case sample complexity[48].
In addition to accuracy, the symbolic models constructed by SISSO and KAN are also worth considering. The symbolic model provided by SISSO is as follows:
$\log{(\kappa)_{{\rm{SISSO}}}}=-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}$ (1)
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to LTC. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to LTC, such as the thermal expansion coefficient[50, 51]. Our feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling LTC.
Table 2: Features adopted in the dataset.
Feature Abbreviation Explanation
space group relaxed $SG$ The space group number for the relaxed structure
AGL Debye temperature $\Theta$ Debye temperature
AGL Grüneisen parameter $\gamma$ Grüneisen parameter
AGL heat capacity constant volume (300K) ${{C}_{\text{v}}}$ Heat capacity per cell at constant volume (300K)
AGL heat capacity constant pressure (300K) ${{C}_{\text{p}}}$ Heat capacity per cell at constant pressure (300K)
AGL thermal expansion coefficient (300K) $\alpha$ Thermal expansion coefficient (300K)
AGL vibrational entropy atom (300K) ${{S}_{\text{v}}}$ Vibrational entropy per atom (300K)
AGL vibrational free energy atom (300K) $F$ Vibrational free energy per atom (300K)
AGL bulk modulus static (300K) ${{K}_{\text{s}}}$ Static bulk modulus (300K)
AGL bulk modulus isothermal (300K) ${{K}_{\text{i}}}$ Isothermal bulk modulus (300K)
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Figure 2(b). The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq.( 1). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert[52]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and LTC, there are a few points to consider:
  • Generally, the symbolic models constructed through machine learning require more physical features (e.g., SISSO involves 7 features, while KAN utilizes all 10 features) compared to semi-empirical models (such as Slack[27], which, when the temperature variable is fixed at 300K, requires only 5 additional features). Moreover, obtaining the corresponding features for unknown materials often involves high experimental or computational costs, making such models appear less appealing at present.
  • In materials inverse design, it is often necessary to regulate dominant features to induce changes in the target physical properties. Although the decision-making process of symbolic models is transparent, relying solely on the symbolic model itself to infer the dominance (or “contribution”) of different features is impractical.
  • Both KAN and SISSO, as white-box models, achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Table S1), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence LTC. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
II.2 Deepening physical interpretability through sensitivity analysis
We identified the most critical features for LTC through two steps. The first step is correlation analysis, achieved by calculating correlation coefficients. The second step is sensitivity analysis, accomplished by computing sensitivity indices.
Figure 3: Feature correlation and sensitivity analysis. (a) A heatmap of correlation coefficients is presented, where the numbers in the upper-right triangle represent the Pearson correlation coefficients between pairs of features. In the lower-left triangle, the direction of the ellipse’s long axis indicates the positive or negative correlation, while the area reflects the strength of correlation. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{v}$) and heat capacity at constant pressure ($C_{p}$), as well as isothermal bulk modulus ($K_{i}$) and static (adiabatic) bulk modulus ($K_{s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Figure 3(a). It is important to clarify that $K_{i}$ and $K_{s}$ are not entirely identical concepts[53, 54, 55]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities[53]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom ${S_{\rm{v}}}$ also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation[56]:
$F=Q-T{S_{\rm{v}}}$ (2)
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300K$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation (EEC)[57] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder[58], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
$H\left({\vec{\sigma},T}\right)={E_{0}}\left({\vec{\sigma}}\right)+\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots$ (3)
[…]
$F\approx\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T$ (4)
where $\left\langle\ln\left(\omega\right)\right\rangle\left(\sigma\right)$ is a constant, thus the vibrational free energy per atom is approximately linearly related to the Debye temperature[58]. These conclusions, combined with Eq.(4) sufficient explanation for the linear relationship between $\Theta$, $F$, and ${{S}_{\text{v}}}$.
We propose that among highly linearly correlated physical features, retaining only one is sufficient. Under ideal conditions, this approach can simplify the ML model without significantly degrading its performance. To achieve this goal, it is necessary to combine correlation analysis with sensitivity analysis. The purpose of sensitivity analysis is to determine which features have a greater impact on the target compared to others, or, more simply, which features are more “important” for the target. Currently, mainstream sensitivity analysis algorithms include SHAP[105], LIME[60], and Sobol[61]. However, these methods generally assume feature independence and the features used in our work are interdependent, their applicability is limited. Ignoring input interactions and multivariate distribution characteristics can severely skew or even invalidate any sensitivity analysis results[62]. Kucherenko et al.[63] improved the traditional Sobol index by using Copula-based sampling to separate each feature’s marginal and joint distributions, thereby constructing a dependency model. In this work, we will implement the Kucherenko indices using UQLab[64] and compare it with other indices.
As shown in Figure 3, we calculated the mean absolute values of various sensitivity indices, which provide the global values for SHAP and LIME. It is important to note that, regardless of the white-box machine learning surrogate model used, results obtained under the assumption of feature independence lack persuasiveness. In Figure 3(b), significant differences exist among the results from different methods, making it difficult to determine which is more reliable. In Figure 3(d), although different analysis methods yield similar feature sensitivities, correlated features should exhibit similar sensitivities[65]. Evidently, the conclusions from Figure 3(b) and (d) do not satisfy this premise. Therefore, conclusions drawn from sensitivity analysis methods that ignore feature interaction effects lack credibility.
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{p}$ and $C_{v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. 1), specific heat capacities appear in the form $\left|{{C_{v}}-{C_{p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that log($\kappa$) is entirely independent of specific heat capacities.
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $log(\kappa)$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs[28, 66]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to log($\kappa$). Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
Figure 4: Performance comparison before and after dimensionality reduction. (a) After extracting key features and re-modeling the LTC, both KAN and the black-box models exhibited good robustness, while the robustness of SISSO did not meet expectations. (b) The simplified architecture of KAN model.
Ultimately, we retained ${{C}{\text{v}}}$, $\alpha$, $F$, and ${{K}{\text{s}}}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq.13), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
After key feature extraction, we performed symbolic regression again using KAN and SISSO. The symbolic model fitted by KAN became considerably simplified, as shown below:
$\log{\left(\kappa\right)_{{\rm{KAN}}}}=$ $-1.05\sin{f_{1}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)+0.61\cosh{f_{2}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)$
$-1.44+1.4{{\rm{e}}^{-19.72f_{3}^{2}\left({F,{K_{s}},\alpha,{C_{v}}}\right)}}-0.48{{\rm{e}}^{-63.21f_{4}^{2}\left({{C_{v}},{K_{s}},F}\right)}}$ (5)
where
${f_{1}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)=$ $-5.66{\left({0.77-F}\right)^{2}}-0.7\tan\left({1.56{K_{s}}-4.4}\right)+2.89$
$+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{v}}-0.11}\right)}^{2}}}}$
${f_{2}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)=$ $-0.76\sin\left({4.23F+4.79}\right)+1.03\tan\left({1.55\alpha-0.99}\right)$
$-0.86\cosh\left({5.35{C_{v}}-1.75}\right)+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{s}}-0.63}\right)+2.31$
${f_{3}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)=$ ${\left({0.38-{K_{s}}}\right)^{3}}-0.09\sin\left({6.2{C_{v}}-6.16}\right)$
$-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)+0.47-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}}$
${f_{4}}\left({{C_{v}},{K_{s}},F}\right)=-{\left({0.37-{C_{v}}}\right)^{2}}-0.13\tan\left({1.46{K_{s}}-0.63}\right)-0.3+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}$
The analytical expression fitted by SISSO remains consistently concise:
[…]
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Figure 4(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between LTC and the features, while also demonstrating exceptional robustness.
II.3 Screening potential thermal insulators/conductors
Unlike simple descriptors that can be directly generated by programs, such as those based on composition (e.g., Magpie[67]) or structure (e.g., Coulomb matrix[68]), the physical features involved in the aforementioned reduced dataset are relatively complex and not easily constructed physical quantities. Inspired by previous studies[69, 70], our approach is to use a two-stage prediction method. We predict features by simple descriptors and models, and then use the predicted features to perform high-throughput predictions of the LTC for unknown materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties[35]. In our application, CGCNNs delivered fairly accurate results for the prediction of all features except for $C_{\text{v}}$, where the prediction accuracy was notably lower (see Supplementary Information for details). Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using features that can be reliably predicted qualitatively, and then calculating $C_{\text{v}}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate LTC.
Figure 5: Correlation between the 4 features and LTC. (a)-(d) The two-dimensional mapping of different features with $\log(\kappa)$, where the color bar reflects the scatter density, and the feature values have been min-max normalized. (e) The mapping between $\log(\kappa)$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa)$ of each sample. The dashed line represents the mean $\log(\kappa)$ value of 1 for the samples.
Figure 6: Using CGCNN to predict key features and qualitatively assess LTC. (a), (b), and (c) display scatter plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets. (d), (e), and (f) illustrate the distribution curves of $F$, $K_{\text{s}}$, and scores for Materials Project samples, ordered by their comprehensive scores, with the red and blue scatter points of varying shapes representing the top 5 and bottom 5 samples, respectively.
Based on the information from Figure 3(a) and Figure 5(a)-(d), it can be concluded that $F$, $K_{\text{s}}$, and $\kappa$ exhibit a strong positive linear correlation. Materials with higher F and $K_{\text{s}}$ values generally exhibit higher LTC, a relationship that is further confirmed in Figure 5(e). Therefore, the preliminary qualitative screening of thermal insulators and conductors based on F and $K_{\text{s}}$ is highly beneficial for accelerating the exploration of new materials. The predictions of F and $K_{\text{s}}$ using CGCNN are shown in Figure 6(a) and 6(b), with $R^{2}$ values reaching 0.94 and 0.98, respectively. We extracted a total of 2,246 samples from the Materials Project[71] with the following selection criteria:
  • Excluding transition metal elements;
  • Number of atoms (nsites) $<$ 5;
  • Band gap $\in\left({0,1.5{\rm{eV}}}\right]$
Additionally, we developed a comprehensive score to qualitatively evaluate a sample’s potential to become a thermal insulator or conductor. This score incorporates the linear coefficients between features and targets, the Kucherenko indices of features with respect to the target, and the features themselves:
${\rm{Score}}=\frac{{{K_{F}}}}{{{K_{F}}+{K_{{K_{s}}}}}}\frac{{{R_{F}}}}{{{R_{F}}+{R_{{K_{s}}}}}}F+\frac{{{K_{{K_{s}}}}}}{{{K_{F}}+{K_{{K_{s}}}}}}\frac{{{R_{{K_{s}}}}}}{{{R_{F}}+{R_{{K_{s}}}}}}{K_{s}}$ (7)
where $K$ represents the respective Kucherenko index, and $R$ denotes the Pearson correlation coefficient. Based on the scores of the materials, we retained the top 5 and bottom 5 samples to verify their potential as thermal conductors or insulators. Existing literature[72, 73, 74, 75, 76, 77] on these samples primarily focuses on 2D materials or electronic thermal conductivity. In contrast, our research targets the LTC of 3D materials, aiming to fill this gap in the field.
Figure 7: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the thermal conductivity to converge. The deep red dashed line represents the thermal conductivity predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The LTC calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) shows a histogram of the thermal conductivity distribution in the original dataset.
Based on the conclusions provided by the MP database, most of the 10 selected structures are predicted to be unstable; consequently, only a few materials, such as CN2, Be2C, CrB4, CuBO2, and BeSiO3, possess valid ${C_{\rm{v}}}$ values obtained through DFT calculations (see Table 3). We utilized GPUMD to calculate both the surface and bulk thermal conductivities of viable candidate materials. For isotropic crystals, we computed results in only one direction. For anisotropic crystals, we calculated the LTC in each direction and averaged these values to represent the overall performance. Notably, CN2 exhibits thermal conductivity exclusively in the $z$-direction. As such, averaging the LTC across multiple directions for this material is not meaningful, and the z-direction thermal conductivity is used to represent its overall performance. Figure 7(a), (b), and (c) illustrate the LTC convergence for each material. We calculated both the surface and bulk thermal conductivities of each material over relaxation times ranging from 0 to 10 ns, providing the sum of these values for comparison with the predictions from the KAN model. In our work, the pretraining model faced several challenges:
  • Feature selection: Initially, the original dataset provided 10-dimensional features for training all models. However, to reduce model complexity, enhance interpretability, and account for the difficulty of obtaining certain features, we applied sensitivity analysis to narrow the features down to just four dimensions. While both KAN and the black-box models demonstrated remarkable robustness, the uncertainty introduced by feature selection remains a consideration.
  • Two-stage prediction: As shown in Figure 6, three features for the candidate materials were derived from CGCNN predictions, with only $C_{\text{v}}$ obtained from DFT calculations. Although CGCNN and DFT provide satisfactory accuracy for individual properties, the two-stage prediction process inevitably introduces cumulative errors in predicting the LTC of candidate materials.
  • Data imbalance[78]: As depicted in Figure 7(d), only 0.47% of the samples in the dataset used for model pretraining have an LTC $>$ 100 W/m/K. Consequently, the model tends to favor the more abundant low- and medium-conductivity materials, potentially limiting its extrapolation capability in unrepresented regions.
According to our MD calculations, the candidate materials exhibit LTC values within distinct ranges (3̃0, 8̃0, 1̃50 W/m/K). Despite this, their ML-predicted LTC consistently align with expectations. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability. Although the qualitative scores for CrB4 and Be2C are quite close, there is a significant difference in their actual thermal conductivities. This discrepancy is attributed to a notable difference in their heat capacities (see Table 3), which is not accounted for in Eq. 6 due to its negligible negative correlation with LTC. However, overall, the qualitative assessment method in this study helped us identify other new materials with excellent thermal conductivity, and it proved valuable in improving the efficiency of material screening.
Table 3: The predicted results of LTC based on KAN.
MPID Formula $F_{pred}$ $K_{s-pred}$ $\gamma_{pred}$ $C_{v-DFT}$ $\kappa_{KAN}$
mp-1009818 CN2 152.8788 333.6628 $1.71^{-5}$ 0.2343 113.5976
mp-27710 CrB4 113.7286 259.3992 $2.73^{-5}$ 12.7181 29.3775
mp-1569 Be2C 138.8222 191.1319 $2.41^{-5}$ 4.6692 80.5562
mp-1096940 CuBO2 78.15269 267.6422 $3.04^{-5}$ 8.4535 31.3254
mp-1183445 BeSiO3 85.25379 224.8873 $2.56^{-5}$ 9.5514 41.3833
mp-1095948 Cs2KNa -139.976 18.569 / / /
mp-1097263 Cs2RbNa -144.966 16.8264 / / /
mp-10378 Cs3Sb -143.5 11.2367 / / /
mp-1097633 Cs2KRb -145.791 14.89677 / / /
mp-635413 Cs3Bi -168.671 13.74844 / / /
III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the LTC of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based LTC modeling has achieved remarkable accuracy, but inprovments in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of LTC modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
In this study, we employed KAN to model LTC, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features, highlighting its superior interpretability in this context. Additionally, the analytical expressions derived from KAN serve as a valuable supplement to traditional empirical models. We also developed a two-stage high-throughput prediction framework, integrating CGCNN to accurately predict the primary physical features required for LTC. For $C_{\text{v}}$, a property that CGCNN cannot predict accurately, DFT calculations were used. Building on this foundation, the predictions from the KAN pretraining model aligned closely with MD validation results, enabling us to successfully identify three new materials, one of which was confirmed as an excellent thermal conductor. Compared to using MD to directly compute LTC for unknown materials, our high-throughput prediction framework requires only DFT calculations of $C_{\text{v}}$ to achieve MD-level predictive accuracy. This work not only accelerates the prediction of material thermal conductivity but also enhances our understanding of the underlying physical mechanisms. By balancing accuracy and interpretability, KAN demonstrates its potential as a powerful tool for advancing materials informatics.
IV METHODS
IV.1 Kolmogorov-Arnold Networks (KANs)
The Kolmogorov-Arnold representation theorem[79] provides that any multivariate continuous function defined on a bounded domain can be expressed as a finite composition of continuous functions of a single variable, combined with the operation of addition[80]. For a differentiable function $f:{{\left[0,1\right]}^{n}}\to\mathbb{R}$:
$f\left(\mathbf{x}\right)=f\left({{x}_{1}},...,{{x}_{n}}\right)=\sum\limits_{q=1}^{2n+1}{{{\Phi}_{q}}\left(\sum\limits_{p=1}^{n}{{{\phi}_{q,p}}\left({{x}_{p}}\right)}\right)}$ (8)
where ${{\phi}_{q,p}}:\left[0,1\right]\to\mathbb{R}$ as well as ${{\Phi}_{q}}:\mathbb{R}\to\mathbb{R}$. In this context, $p$ denotes the number of top operators, $q$ denotes the number of bottom operators, and $n$ denotes the number of nodes in the bottom network (which corresponds to the dimensionality of the input features in the input layer), $\mathbf{x}=\left({{x}_{1}},...,{{x}_{n}}\right)$ represents feature vector.
In the original K-A representation theorem, the number of nonlinear layers is limited to 2, and the number of hidden layer nodes is set to $12n+$ 1, which fixes the network structure to $\left[n,2n+1,1\right]$. However, Liu et al.[30] are not constrained by these limitations. In KANs, both the number of layers and the width of the network are arbitrary, which enhances the feasibility of the K-A representation theorem for ML applications. Therefore, for a KAN with ${{n}_{0}}$-D inputs, L layers, and ${{n}_{\text{L}}}=1$-D output, a more precise definition is:
$\text{KAN}\left(\mathbf{x}\right)=\sum\limits_{{{i}_{L-1}}=1}^{{{n}_{L-1}}}{{{\phi}_{L-1,{{i}_{L}},{{i}_{L-1}}}}\left(\sum\limits_{{{i}_{L-2}}=1}^{{{n}_{L-2}}}{\cdots\left(\sum\limits_{{{i}_{2}}=1}^{{{n}_{2}}}{{{\phi}_{2,{{i}_{3}},{{i}_{2}}}}\left(\sum\limits_{{{i}_{1}}=1}^{{{n}_{1}}}{{{\phi}_{1,{{i}_{2}},{{i}_{1}}}}\sum\limits_{{{i}_{0}}=1}^{{{n}_{0}}}{{{\phi}_{0,{{i}_{1}},{{i}_{0}}}}\left({{x}_{{{i}_{0}}}}\right)}}\right)}\right)\cdots}\right)}$ (9)
the expression can be simplified to:
$\text{KAN}\left(\mathbf{x}\right)=\left({{\mathbf{\Phi}}_{L-1}}\circ{{\mathbf{\Phi}}_{L-2}}\circ\cdots\circ{{\mathbf{\Phi}}_{1}}\circ{{\mathbf{\Phi}}_{0}}\right)\mathbf{x}$ (10)
in the expression, ${{\mathbf{\Phi}}_{i}}=\left\{{{\phi}_{q,p}}\right\}$ denotes the i-th layer of the KAN, defined as a tensor of 1-D activation functions ${{\phi}_{q,p}}$. In the implementation details, $\phi\left(x\right)$ is represented as a linear combination of the basis function $b\left(x\right)$ and B-spline function:
$\phi\left(x\right)=\omega\left[b\left(x\right)+\text{spline}\left(x\right)\right]=\omega\left[\frac{x}{1+{{\text{e}}^{-x}}}+\sum\limits_{i}{{{c}_{i}}{{B}_{i}}\left(x\right)}\right]$ (11)
where B-spline function is defined as:
$B_{i,0}(x)$ =$=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1},\\ 0,&\text{otherwise}.\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x).$ (12)
In KANs, the coefficients $c$, the degree $k$, and the grid $G$ of the spline functions are all learnable parameters, which are updated through backpropagation.
IV.2 Sure Independence Screening & Sparsifying Operator (SISSO)
SISSO constructs a descriptor vector ${{\mathbf{d}}_{n}}$ using features and operators, and models the target vector $\mathbf{P}$ through a linear combination of $n$-dimensional descriptors. The initial feature ${{\mathbf{\Phi}}_{0}}$ serves as the starting point for constructing descriptors and includes easily obtainable physical parameters. The operator set is defined as:
${{\hat{H}}^{\left(\text{m}\right)}}\equiv\left\{I,+,-,\times,\div,\exp,\log,\text{abs},\sqrt{{}}{{,}^{-1}}{{,}^{-2}}{{,}^{-3}}\right\}\left[{{\phi}_{1}},{{\phi}_{2}}\right]$ (13)
where ${{\phi}_{1}}$ as well as ${{\phi}_{2}}$ are terms in ${{\mathbf{\Phi}}_{i}}$. The superscript ${}^{\left(\text{m}\right)}$ indicates that SISSO retains only descriptors with physical meaning. For example, features that are added or subtracted must have the same dimensions, and features involved in logarithmic or square root operations cannot be negative. The new features constructed during the i-th iteration can be expressed as
${{\Phi}_{i}}=\bigcup\limits_{k}{{{{\hat{h}}}_{k}}^{\left(\text{m}\right)}\left[{{{\hat{\phi}}}_{i}},{{{\hat{\phi}}}_{j}}\right]},\forall{{\hat{h}}_{k}}^{\left(\text{m}\right)}\in{{\hat{H}}^{\left(\text{m}\right)}}\text{ and }\forall{{\phi}_{i}},{{\phi}_{j}}\in{{\Phi}_{i-1}}$ (14)
where ${{\hat{h}}_{k}}^{\left(\text{m}\right)}$ represents single operator of ${{\hat{H}}^{\left(\text{m}\right)}}$, ${{\phi}_{i}}$ and ${{\phi}_{j}}$ are different elements from ${{\Phi}_{i-1}}$. SIS[81] constructs the feature subspace with the highest correlation to the target $\mathbf{P}$ through vector inner products, while SO selects the top n features with the highest relevance using regularization techniques to form the descriptor matrix ${{\mathbf{d}}_{n}}$. The linear coefficients can be approximately determined by solving the equation ${{\mathbf{c}}_{n}}={{\left(\mathbf{d}_{n}^{\text{T}}{{\mathbf{d}}_{n}}\right)}^{-1}}\mathbf{d}_{n}^{\text{T}}\mathbf{P}$, so the model constructed by SISSO can be represented as:
$\mathbf{\hat{P}}={{\mathbf{d}}_{n}}{{\mathbf{c}}_{n}}={{c}_{0}}+{{c}_{1}}{{d}_{1}}+\cdots+{{c}_{n}}{{d}_{n}}$ (15)
where $\mathbf{\hat{P}}$ represents the estimated target vector of the model.
IV.3 Sobol Indices
Sobol is a method for feature importance analysis based on analysis of variance (ANOVA)[61]. It allocates a portion of the total variance to each input variable or its interactions with other variables, thereby providing valuable information about the importance of each input variable[66]. The first-order and total Sobol indices can be defined as follows:
${{S}_{i}}=\frac{\text{Var}\left[{{\text{E}}_{\sim i}}\left[Y|{{X}_{i}}\right]\right]}{\text{Var}\left[Y\right]}$ (16)
$S_{i}^{\text{T}}=\frac{{{\text{E}}_{\sim i}}\left[\text{Var}\left[Y|{{\mathbf{X}}_{\sim i}}\right]\right]}{\text{Var}\left[Y\right]}$ (17)
for a specific value of ${{X}_{i}}$, the value of $Y$ can be determined by averaging the model evaluations over a sample of ${{\mathbf{X}}_{\sim i}}$ while keeping ${{X}_{i}}=x_{i}^{*}$ fixed, where ${{\mathbf{X}}_{\sim i}}$ represents all variables except ${{X}_{i}}$.
IV.4 Kucherenko Indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies[63, 66, 82]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas[83] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure[84], as detailed in Table S2. This approach approximates the correlated variables as independent before calculating the Sobol indices. The Gaussian copula is constructed based on the standard normal distribution, and its mathematical expression is:
$C\left({{u_{1}},{u_{2}},\ldots,{u_{d}};\Sigma}\right)={\Phi_{\Sigma}}\left[{{\Phi^{-1}}\left({{u_{1}}}\right),{\Phi^{-1}}\left({{u_{2}}}\right)\ldots,{\Phi^{-1}}\left({{u_{d}}}\right)}\right]$ (18)
where $u_{i}=F_{i}(x_{i})$ is the Cumulative Distribution Function (CDF) of the marginal distribution, $\Phi^{-1}$ is the inverse of the CDF of the standard normal distribution $\mathcal{N}(0,1)$, $\Phi_{\Sigma}$ is the CDF of a multivariate normal distribution with mean zero and covariance matrix $\Sigma$, and $\Sigma$ is the correlation matrix, describing the dependence structure between the random variables.
IV.5 Dataset construction & feature preprocessing
The dataset used in this work is sourced from the aflowlib.org database[85], comprising a total of 5,578 entries. For feature selection, the focus was primarily on characteristics related to vibrational, thermodynamic, and mechanical properties at room temperature (300K), which are theoretically associated with LTC according to first-principles researches[86, 87]. For more details, please refer to Table 2. In the modeling aimed at LTC, the model predicts the logarithm of $\kappa$ (log($\kappa$)) rather than the LTC itself. This is because the logarithmic transformation compresses the target value space, thereby enhancing the performance of the ML models[88]. For deep models such as MLPs and KANs, normalization of features is essential as these models rely on gradient descent for parameter optimization. Normalization enhances numerical stability by ensuring that features with different ranges contribute equally to the gradient computations; otherwise, features with larger values might dominate the optimization process[89]. However, this step is unnecessary for XGBoost[90], as it is based on decision trees. Additionally, for interpretability reasons, feature normalization is often omitted in recent works based on SISSO[26, 28, 91]. In this work, we employed Min-Max Normalization, which is expressed as:
$x^{\prime}=\frac{{x-{x_{\min}}}}{{{x_{\max}}-{x_{\min}}}}$ (19)
where $x$ is the original feature value, ${{x}_{\min}}$ and ${{x}_{\max}}$ are the minimum and maximum values of the feature, respectively, and ${x}^{\prime}$ is the normalized feature value.
IV.6 Implementation of DFT & GPUMD methods
DFT and AIMD calculations are performed using VASP[92, 93]. The Projector-Augmented Wave (PAW) method[94] and the Perdew-Burke-Ernzerhof (PBE) functional[95] within the Generalized Gradient Approximation (GGA) are employed for electron exchange-correlation. The cutoff energy is set to 500 eV, and the electronic convergence threshold is $10^{-8}$ eV.
For the calculation of LTC, both the phonon Boltzmann Transport Equation (BTE) method and Molecular Dynamics (MD) simulations are used. For the stable Be2C structure, specific heat and LTC are computed using ShengBTE[96]. The force constants are generated with Hiphive[97]. For other structures, specific heat is calculated using Phonopy[98], while LTC is determined through GPUMD[99]. AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ps. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,\text{\AA}^{-1}$ has been tested and found to be optimal for this study.
IV.7 Data Availability
The original LTC data and the selected features can be downloaded from https://aflowlib.org/ or accessed via the GitHub homepage at https://github.com/FlorianTseng/LTC-modeling.
IV.8 Code Availability
The codes supporting our research are available at https://github.com/FlorianTseng/LTC-modeling.
IV.9 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements.
We would like to acknowledge the National Key R&D Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12122408, No.12404062, No.12302220, No.1237041184), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO. ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No. 2042023kf0109). We sincerely thank the High-Preformance Computing Center of Wuhan University for providing computational resources for our work.
Supplementary Information
Supplementary Note 1 KAN’s symbolic model on complete dataset
The simplified symbolic model in the main text Eq.(5), derived from key physical features, is sufficient for an accurate description of LTC. However, for reference, we also provide the symbolic expression fitted by KAN using the complete dataset. It is worth noting that, due to the architecture of KAN, this formula is inevitably more complex. Therefore, we present it solely in its Python representation:
1 log(kappa) =
[…]
Supplementary Note 2 Multi-Layer Perceptrons (MLPs)
The theoretical framework of MLPs is well-established. Therefore, in this work, we emphasize the distinctions between MLPs and KANs. Unlike KANs, MLPs are founded on the universal approximation theorem[100], which can be mathematically expressed as:
$f\left(x\right)\approx\sum\limits_{i=1}^{N}{{{\alpha}_{i}}\sigma\left(\omega_{i}^{\text{T}}x+{{b}_{i}}\right)}$ (Supplementary Eq.1)
where $\alpha$ are the coefficients, $\sigma$ is the activation function (we adopted ReLU[101] in this work), ${{\omega}_{i}}$ and ${{b}_{i}}$ denote the weights and biases, $N$ represents the number of neurons in the hidden layer. The expression of MLPs can be simplified to:
$\text{MLP}\left(\mathbf{x}\right)=\left({{\mathbf{W}}_{L-1}}\circ\sigma\circ{{\mathbf{W}}_{L-1}}\circ\sigma\circ\cdots\circ{{\mathbf{W}}_{1}}\circ\sigma\circ{{\mathbf{W}}_{0}}\right)\mathbf{x}$ (Supplementary Eq.2)
In MLPs, only the weights and biases, which serve as linear parameters, are learnable, while the nonlinear activation functions are fixed as hyperparameters.
Supplementary Note 3 eXtreme Gradient Boost (XGB)
XGB represents a great advancement over the traditional Gradient Boosting Decision Tree (GBDT) algorithm. Its innovations include the incorporation of regularization to enhance generalization, a two-step gradient approximation of the objective function to accelerate computational efficiency, column subsampling to mitigate noise and further bolster generalization, and an advanced method for handling missing values by assigning a default classification direction at tree nodes, which renders it particularly effective for sparse datasets[102]. GBDT algorithms, including XGB, are computationally intensive particularly in tasks with high-dimensional features and large datasets[102, 103].
Supplementary Note 4 Crystal Graph Convolutional Neural Network (CGCNN)
CGCNN[104] directly learns from the connection of atoms in a crystal, representing the crystal as a graph with atoms as nodes and bonds as edges. This graph-based approach enables the model to automatically extract optimal representations for predicting material properties. The key feature of CGCNN is its Convolutional Neural Network (CNN) built on top of these crystal graphs. The convolutional layers update the feature vectors of atoms by incorporating information from neighboring atoms and bonds, using the following graph convolution function:
$v_{i}^{\left(t+1\right)}=\text{Conv}\left(v_{i}^{\left(t\right)},v_{j}^{\left(t\right)},u_{\left(i,j\right)}^{\left(k\right)}\right),{{\left(i,j\right)}_{k}}\in G$ (Supplementary Eq.3)
where $v_{i}^{\left(t\right)}$ is the feature vector of atom i at layer t, and $u_{\left(i,j\right)}^{\left(k\right)}$ is the feature vector of the bond between atoms i and j. After multiple convolutional layers, the pooling layers aggregate the atomic feature vectors into an overall crystal representation, which is then passed to fully connected layers to predict material properties.
Supplementary Note 5 SHAP Indices
Shapley Additive Explanations (SHAP)[105] is derived from the concept of Shapley values[106] in game theory, which were originally employed to assess the contribution of each individual player to the collective gain within a coalition[107]. Shapley value for individual player can be expressed as:
${{\phi}_{i}}=\sum\limits_{S\subseteq N\backslash\left\{i\right\}}{\frac{\left|S\right|!\left(\left|N\right|-\left|S\right|-1\right)!}{\left|N\right|!}}\left(v\left(S\cup\left\{i\right\}\right)-v\left(s\right)\right),i=1,2,\ldots,N$ (Supplementary Eq.4)
where $N$ represents the set of all features, $S$ is any subset of features that does not include feature $i$, $\left|S\right|$ is the number of features in set $S$, is the contribution of feature set $S$ to the model’s prediction output, and
$v\left(S\cup\left\{i\right\}\right)$
is the contribution of the feature set
$S\cup\left\{i\right\}$
, which includes feature $i$, to the model’s prediction output. The core concept of SHAP is to generate explanations for individual predictions, which can then be aggregated to provide both local and global insights into the model’s behavior[105], it can be expressed mathematically as:
$\hat{f}\left(\mathbf{x}\right)={{\phi}_{0}}+\sum\limits_{i=1}^{N}{{{\phi}_{i}}\left(\mathbf{x}\right)}$ (Supplementary Eq.5)
where
${{\phi}_{i}}\left(\mathbf{x}\right)$
is the Shapley value of the i-th variable, ${{\phi}_{0}}$ is typically set as the expected value of the training samples, $\hat{f}\left(\mathbf{x}\right)$ represents the original prediction made by the model[108].
Supplementary Note 6 LIME Indices
Local Interpretable Model-agnostic Explanations (LIME)[60] is a perturbation-based method for feature importance analysis that employs a local surrogate model to interpret individual samples. By generating perturbations around the selected sample, a new dataset is created, and an interpretable model (such as linear regression) is trained. The difference between the two models is then measured using an objective function:
$\xi\left(x\right)=\arg\underset{g\in G}{\mathop{\min}}\,L\left(f,g,{{\pi}_{x}}\right)+\Omega\left(g\right)$ (Supplementary Eq.6)
in the expression, $f$ denotes the model to be explained, $g$ represents a simplified model drawn from a set $G$ of potential models (such as linear models), ${{\pi}_{x}}$ quantifies the distance between new samples ${x}^{\prime}$ and the original samples $x$, and $\Omega\left(g\right)$ reflects the complexity of the model $g$.
Supplementary Note 7 Supplementary Tables
Supplementary Table 1: The optimal hyperparameters.
Model Hyperparameter Optimal Value
MLP hidden_sizes [64, 64, 64] (fixed)
learning_rate 0.002587427
batch_size 186
num_epochs 592
XGB n_estimators 489
learning_rate 0.134952344
max_depth 6
subsample 0.824230319
colsample_bytree 0.869698424
reg_alpha 0.9
reg_lambda 37.30538969
KAN width 7
grid 15
k 5
lamb_l1 9.559785284
steps 27
lr 0.47852298
SISSO (10-D) desc_dim 2
nsf 10
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 5
funit (1:10)
nmodels 100
nf_sis 10
SISSO (4-D) desc_dim 2
nsf 4
ops (+)(-)(*)(/)(exp)(exp-)(^-1)(^2)(^3)(sqrt)(cbrt)(log)(—-—)(scd)(^6)(sin)(cos)
fcomplexity 4
funit (1:4)
nmodels 100
nf_sis 10
Supplementary Table 2: Marginal distribution of individual primary features determined by Copula sampling.
Distribution Parameters
$\Theta$ lognorm $\mu=-1.935,\sigma=0.519$
$\gamma$ lognorm $\mu=1.671,\sigma=0.012$
$C_{p}$ lognorm $\mu=-2.611,\sigma=0.778$
$C_{v}$ lognorm $\mu=-2.674,\sigma=0.831$
$\alpha$ lognorm $\mu=-2.370,\sigma=0.507$
$S_{v}$ norm $\mu=0.323,\sigma=0.119$
$F$ lognorm $\mu=-0.027,\sigma=0.102$
$K_{s}$ gamma $k=2.546,\theta=0.102$
$K_{i}$ gamma $k=2.441,\theta=0.104$
Supplementary Note 8 Supplementary Figure
Supplementary Figure 1: Histogram of the marginal distribution of a single key variable.
The variables involved in Table S2 are classified into three types of distributions: Gamma distribution, normal distribution, and log-normal distribution. Their respective probability density functions (PDF) can be expressed as:
$f\left(x\right)$ $=\frac{1}{{\Gamma\left(k\right){\theta^{k}}}}{x^{k-1}}{{\rm{e}}^{-\frac{x}{\theta}}},$ (Supplementary Eq.7)
$f(x)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right),$ (Supplementary Eq.8)
$f\left(x\right)$ $=\frac{1}{{x\sigma\sqrt{2\pi}}}{{\rm{e}}^{-\frac{{{{\left[{\ln\left(x\right)-\mu}\right]}^{2}}}}{{2{\sigma^{2}}}}}}.$ (Supplementary Eq.9)